Calculate the number of grams of barium chloride dehydrate required to react completely with 6.77 grams of potassium sulfate.
1
Expert's answer
2017-11-15T12:55:07-0500
K2SO4 + BaCl2*2H2O = BaSO4 + 2 KCl + 2 H2O
n (K2SO4) = m/M = (6.77 g)/(174 g/mol) = 0.039 mol
n (BaCl2*2H2O) = n (K2SO4) = 0.039 mol
m (BaCl2*2H2O) = n * M = 244 g/mol * 0.039 mol = 9.516 g
Answer:
m (BaCl2*2H2O) = 9.516 g